# Design-Specific Transforms In Visualization

### v1

@@ 3.3 Data Proxies and Analytic Shortcuts
The preceding examples have assessed the effectiveness of a visualization’s pre-computation by describing the steps involved to answer the task from the perspective of the data-level proxy query (data proxy). The underlying implication is that the difficulty of using the visualization is proportional to the complexity of the data proxy. For instance, parts-to-whole calculations are a simple read operation using a pie chart, but require mental arithmetic when using a scatterplot or bar chart. Since the data proxy can be described as a sequence of primitive operations (e.g., filter, read, calculate, etc), it may be possible to assign costs to each operation to estimate the difficulty of the data proxy, and consequently to the view-level proxy query. Such a “cost model” is analogous to those used to evaluate user interfaces [11] and SQL queries [25], and could e.g., serve as a null model for evaluating visualization task-effectiveness, rank visualization designs, or evaluate the difficulty of multi-step tasks.
Of course, a key drawback of this idea is that users are not computers—they make use of visual and cognitive heuristics, as well as domain expertise and past experience. How can this be adequately modeled in a way that is useful?
Consider the task of estimating the slope of the regression line fit to a set of points in Figure 2. The left side pre-computes the task by fitting and rendering a line so the user can directly read the slope. In contrast, expert visualization users may develop analytic shortcuts, such as the use of visual features as proxies to solving the complex mental arithmetic[30, 13] (right side). Thus, although the scatterplot does not pre-computate, the task is easy to perform by using such shortcuts [4, 5].
One way to model this is by extending the set of primitive operations with a library meta-operations that are cheap for the user to evaluate, but covers multiple steps in the data-level proxy query $\tilde{q}^{P}$. Different expertise levels can be modeled as different libraries of meta-operators.
@@
3.4 Attribution-Based Experiment Design
When evaluating different visualization designs, we should consider design-specific transformations and the visual encoding as separate experimental factors. As such, two visual mappings only make sense to compare if they differ in only one of the factors, or if all combinations are evaluated.
Evaluating Mental Arithmetic. Consider the task “read the value of a bar” using a bar chart and a proportional stacked bar chart. For a bar chart, the visualization encodes the desired information, which a viewer can readily retrieve. For a proportional stacked bar chart, assuming that the user knows the sum ($sum(s)$), the user can inverse the mapping function from proportions information back to absolute values. In this sense, we note that it is appropriate to compare a bar chart and a stacked bar chart for the given task. However, since the two visualizations use similar visualization elements, a comparison between a bar chart and a proportional stacked bar chart is effectively measuring the cost of the mental arithmetic to invert the sum operation.
Evaluating Visual Design. Consider the task “compare two values” using a pie chart and a proportional stacked bar chart. The two visualizations require the same preparatory computations and visualize the same resulting derived information. However, the two differ in their visual encodings, with a pie chart using wedges in a Polar coordinate and the proportional stacked bar chart using rectangles in a Cartesian coordinate system. In this sense, we note that it is appropriate to compare a pie chart and a stacked bar chart for the given task. However, since the preparatory computations and the derived information are the same, the comparison between the two visualizations is effectively measuring the relative perceptual effectiveness of their visualization encodings.
@@ 4 Conclusions and Discussion
In this paper, we present a transformation-centric model that extends the Information Visualization Reference Model (Infovis Reference Model) by Card et al. [7]. Our model separates design-specific transformations from visual encodings, offering two theoretical framings—a view-level proxy query and a data-level proxy query—for considering how transformed data and the resulting visualizations can be used to address a user’s data tasks. Based on our model, we propose the “No Free Lunch” conjecture for visualization design and a framework based on the tradeoff between pre-computation and visualization task flexibility. Finally, we argue that this model provides clearer insights into the relationships between different visualization designs based on a theoretical cost model and informs better experiment design by isolating visual encoding and transformations.

### v2

@@ 4.2 Data Proxies and Analytic Shortcuts
The preceding examples have assessed the effectiveness of a visualization’s pre-computation by describing the steps involved in answering the task from the perspective of the data-level proxy query (data proxy). The underlying implication is that the difficulty of using the visualization is proportional to the complexity of the data proxy. For instance, parts-to-whole calculations are a simple read operation using a pie chart but require mental arithmetic when using a scatterplot or bar chart. Since the data proxy can be described as a sequence of primitive operations (e.g., filter, read, calculate, etc.), it may be possible to assign costs to each operation to estimate the difficulty of the data proxy, and consequently to the view-level proxy query. Such a “cost model” is analogous to those used to evaluate user interfaces [18] and SQL queries [35], and could e.g., serve as a null model for evaluating visualization task-effectiveness, rank visualization designs, or evaluate the difficulty of multi-step tasks.
Of course, a key drawback of this idea is that users are not computers—they make use of visual and cognitive heuristics, as well as domain expertise and experience. How can this be adequately modeled in a way that is useful?
Consider the task of estimating the slope of the regression line fit to a set of points in Figure 7. The left side pre-computes the task by fitting and rendering a line so the user can directly read the slope. In contrast, expert visualization users may develop analytic shortcuts, such as the use of visual features as proxies to solving the complex mental arithmetic[43, 21] (right side). Thus, although the scatterplot does not require pre-computation, the task is easy to perform by using such shortcuts [7, 8].
One way to model this is by extending the set of primitive operations with a library of meta-operations that are cheap for the user to evaluate and cover multiple steps in the data-level proxy query $\tilde{q}^{P}$. Different expertise levels can be modeled as different libraries of meta-operators.
Application to Experiment Design. We now sketch a potential way where a cost model can be used to derive a baseline when evaluating visualization effectiveness for specific tasks. Given a task, we can use the cost model to find the lowest cost strategy, and use that as a “null model” of the user’s expected strategy for answering the view proxy. If the results of a user study coincide with the null model, then it suggests that the user is using the lowest cost strategy i.e., a rational strategy. If the results are worse than predicted by the null model, then it suggests that the user is performing an inefficient strategy, perhaps due to the visual design. Again, it may be possible to enumerate a handful of higher-cost strategies to design follow-up experiments to isolate the specific strategy that user uses. Finally, if the results are better than predicted, it suggests that the user is performing analytic shortcuts that can be further investigated.
@@ 5 Conclusions and Discussion
In this paper, we present a transformation-centric model that extends the Information Visualization Reference Model (InfoVis Reference Model) by Card et al. [10]. Our model separates design-specific transformations from visual encodings, offering two theoretical framings—a view-level proxy query and a data-level proxy query—for considering how transformed data and the resulting visualizations can be used to address a user’s data tasks. Based on our model, we propose the “No Free Lunch” conjecture for visualization design and a framework based on the trade-off between pre-computation and visualization task flexibility. Finally, we argue that this model provides clearer insights into the relationships between different visualization designs based on a theoretical cost model and informs better experiment design by isolating visual encoding and transformations.